OpenStudy (anonymous):

The population of a village is 5000. It is increasing at the rate of 10% each year. At the end of 3 years, 1155 people migrate from the village. The net increase in population at the end of 3 years is:

OpenStudy (lgbasallote):

500 is not constant @TheViper

OpenStudy (theviper):

omg right

OpenStudy (theviper):

1year=5000+500=5500 2year=5500+550=6050 3year=6050+605=??

OpenStudy (lgbasallote):

hmm....you actually use geometric sequence here \[\Large a_n = a_1r^{n-1}\] where a_n will be the population after 3 years <--what you're looking for a_1 will be 5000 because that's the starting population. r will be 1.10 because it increases by 10% each year so 1 + 0.10 = 1.10 n will be 3 because you're looking for 3 years \[\Large \implies a_n = (5000)(1.10)^{3-1}\] tehn to get the final answer just subtract 1155 \[\Large \implies (5,000)(1.10)^{3-1} - 1,155\]

OpenStudy (lgbasallote):

i wonder where @selvi is